# Evaluating the Evolution of YOLO (You Only Look Once) Models: A Comprehensive Benchmark Study of YOLO11 and Its Predecessors

### v1

@@ IV-B4 Overall Performance
Considering accuracy, speed, size, and GFLOPs, YOLO11m, YOLOv11n, YOLO11s, and YOLOv10s emerged as the most consistent performers. They achieved high accuracy, low processing time and power, and efficient disk usage, making them suitable for a wide range of applications where both speed and accuracy are crucial.
Conversely, YOLOv9e, YOLOv5ux, and YOLOv3u demonstrated poor results across all metrics, being computationally inefficient and underperforming relative to their sizes. YOLO11 models showed the best overall performance, likely due to recent enhancements such as the C3k2 block and C2PSA module. Following closely, YOLOv10 models, despite slightly underperforming in accuracy excelled in efficiency thanks to its use of implementation of One-to-One head for prediction. While YOLOv9 showed underperformance in computational efficiency, it remains competitive with YOLOv10 and YOLO11 in terms of accuracy, thanks to its PGI integration. This positions YOLOv9 as a viable choice for applications where precision is prioritized over speed.
@@
In addition, YOLOv8 and YOLOv5u exhibited competitive results, surpassing YOLOv3u in accuracy, which is likely due to YOLOv3u’s older architecture. However, their accuracy still fell significantly short compared to the newer models, such as YOLOv9, YOLOv10, and YOLO11. While YOLOv8 and YOLOv5u had faster processing times than YOLOv9, their overall performance remains inferior to that of the newer models.
IV-B5 Object Size and Rotation Detection
The YOLO algorithm is effective in detecting large and medium-sized objects, as evidenced by high accuracy in the Africa Wildlife and Traffic Signs datasets. However, it struggles with small object detection, probably due to its division of images into grids, making identifying small, low-resolution objects challenging. Adding to that, YOLO faces challenges when handling objects of different rotations due to the inability to enclose rotated objects, leading to sub-par results overall.
To handle rotated objects, models such as YOLO11 OBB [26] and YOLOv8 OBB [25] (Oriented Bounding Box) can be implemented. Keeping the same foundational architecture as the standard YOLOv8 and YOLO11, YOLOv8 OBB and YOLO11 OBB replace the standard bounding box prediction head with one that predicts the four corner points of a rotated rectangle, allowing for more accurate localization and representation of arbitrarily oriented objects.
IV-B6 The Rise of YOLO11 Over YOLOv8
@@
IV-B7 Dataset Size
The size of the dataset significantly influences the performance of YOLO models. For instance, large models did not perform optimally on the small African wildlife dataset compared to their results on the Traffic Signs and Ships and Vessels datasets due to being more prone to overfitting. Conversely, small models such as YOLOv9t and YOLOv9s performed significantly better on the Africa Wildlife dataset compared to their results on the other datasets, showcasing the effectiveness of small-scaled models when handling limited datasets.
IV-B8 Impact of Training Datasets
@@ V Conclusion
How do different YOLO versions perform on datasets with diverse object characteristics, such as size, aspect ratio, and density?
What are the specific strengths and limitations of each YOLO version, and how can these insights inform the selection of the most suitable algorithm for various applications?
In particular, the YOLO11 family emerged as the most consistent, with YOLO11m striking an optimal balance between accuracy, efficiency, and model size. While YOLOv10 delivered slightly lower accuracy than YOLO11, it excelled in speed and efficiency, making it a strong choice for applications requiring efficiency and fast processing. Additionally, YOLOv9 performed well overall and particularly stood out in smaller datasets. These findings provide valuable insights for industry and academia, guiding the selection of the most suitable YOLO algorithms and informing future developments and enhancements. While the evaluated algorithms demonstrate promising performance, there is still room for refinement. Future research could focus on optimizing YOLOv10 to enhance its accuracy while preserving its speed and efficiency advantage. Additionally, continued advancements in architectural design may pave the way for even more groundbreaking YOLO algorithms. Our future work includes in-depth studies of the identified gaps in these algorithms, along with proposed improvements to demonstrate their potential impact on overall efficiency.

### v2

@@ IV-B4 Overall Performance
Considering accuracy, speed, size, and GFLOPs, YOLO11m, YOLO11n, YOLO11s, and YOLOv10s emerged as the most consistent performers. These models demonstrated a strong balance between high accuracy, low processing time, and efficient disk usage, making them suitable for a wide range of applications where both speed and accuracy are crucial. YOLO11 models, in particular, benefited from recent enhancements such as the C3k2 block and C2PSA module, which contributed to their superior performance in these areas. The incorporation of these innovations allowed the YOLO11 models to handle larger datasets more effectively without compromising on speed or efficiency.
On the other hand, YOLOv9e, YOLOv5ux, and YOLOv3u showed poor results across all metrics. These models were computationally inefficient, struggling with high GFLOPs counts and slower processing times. YOLOv3u, in particular, ranked poorly due to its reliance on the outdated Darknet-53 architecture, which is no longer optimized for modern object detection tasks. Despite its larger size, it failed to offer the necessary accuracy, making it less suitable for practical applications. Similarly, YOLOv5ux and YOLOv9e showed poor performance in terms of both speed and accuracy, which further limits their real-world usability.
YOLO11 models were clearly the best performers overall, owing to their modern design and architectural improvements, making them highly suitable for a wide range of use cases. YOLOv10 models, although slightly underperforming in accuracy, excelled in computational efficiency due to their use of the One-to-One head for predictions. This design allowed YOLOv10 to process data more quickly and efficiently, making it ideal for time-sensitive applications where speed is prioritized over perfect accuracy.
While YOLOv9 models showed underperformance in terms of computational efficiency, they still remain competitive with YOLOv10 and YOLO11 in terms of accuracy, particularly due to their PGI integration, which helps in enhancing detection precision. This makes YOLOv9 a viable option for applications where high accuracy is prioritized over speed.
Concerning the YOLOv12 family, it showcased mediocre results, underperforming compared to its predecessors, particularly YOLOv9 and YOLO11, in both accuracy and speed. Despite introducing innovative components like the Area Attention Module (A2), Residual Efficient Layer Aggregation Networks (R-ELAN), and FlashAttention, these architectural changes contributed to increased computational overhead and slower inference times. The Area Attention Module segmented the feature map to maintain a large receptive field but limited contextual understanding, leading to lower accuracy. Meanwhile, R-ELAN improved feature aggregation but introduced latency due to its complex residual connections. FlashAttention aimed to optimize memory access but added processing overhead, further slowing down the model. Although YOLOv12 leveraged convolution operators for efficiency, it failed to offset the complexity introduced by its attention mechanisms. Overall, YOLOv12 struggled to balance speed and accuracy, making it less favorable compared to the more efficient and accurate YOLOv9, YOLOv10, and YOLO11 models.
@@
In addition, YOLOv8 and YOLOv5u achieved competitive results, surpassing YOLOv3u in terms of accuracy. This improvement can be attributed to their more recent architectures, which allowed them to handle modern detection tasks better. However, despite their higher accuracy compared to YOLOv3u, they still fell short when compared to newer models like YOLOv9, YOLOv10, and YOLO11 in both accuracy and overall performance. While YOLOv8 and YOLOv5u had faster processing times than YOLOv9, their overall performance was still inferior to that of the YOLO11 and YOLOv10 models, which maintained a better balance between speed and accuracy.
IV-B5 The Rise of YOLO11 Over YOLOv8
@@
This paper contributes by providing empirical evidence of YOLO11’s superior performance in both accuracy and computational efficiency. Our findings establish YOLO11 as a clear advancement in the YOLO family, setting a new benchmark for object detection. By incorporating novel modules, YOLO11 addresses both the need for high accuracy and low-latency processing, offering valuable insights for deploying YOLO-based models in practical, real-world applications.
IV-B6 YOLOv9: Efficient Yet Slow
Although YOLOv9 [70] excels in GFLOPs count and model size, outperforming many other models, it paradoxically ranks as one of the slowest in terms of inference speed. This unusual combination of high computational efficiency but low speed can be attributed to YOLOv9’s architectural choices, such as the use of Programmable Gradient Information (PGI) and Generalized Efficient Layer Aggregation Network (GELAN), which optimize computational performance but add significant latency during inference.
This paper contributes by highlighting the trade-offs in YOLOv9’s design, showing that while its smaller model size and lower GFLOPs count make it ideal for resource-constrained environments, its slow inference times make it unsuitable for time-sensitive applications. Our findings suggest that YOLOv9’s performance is best suited for scenarios where computational efficiency is prioritized over speed, offering valuable insights for selecting models based on specific application needs.
IV-B7 YOLOv12: Ambitious but Underwhelming
YOLOv12, despite being the latest addition to the YOLO family, delivered subpar results, failing to outperform its predecessors, particularly YOLOv9, YOLOv10, and YOLO11, in both accuracy and speed. This underperformance is primarily due to its complex architectural choices, including the Area Attention Module (A2), Residual Efficient Layer Aggregation Networks (R-ELAN), and FlashAttention, which introduced significant computational overhead. Although these modules aimed to enhance contextual understanding and memory efficiency, they ultimately compromised speed and accuracy due to increased latency and reduced spatial awareness. This work contributes by revealing the trade-offs inherent in YOLOv12’s ambitious design, demonstrating that its added complexity did not translate into superior performance.
IV-B8 Dataset Size
The size of the dataset significantly influences the performance of YOLO models. For instance, large models did not perform optimally on the small African wildlife dataset compared to their results on the Traffic Signs and Ships and Vessels datasets due to being more prone to overfitting. Conversely, small models like YOLOv9t and YOLOv9s performed best on the Africa Wildlife dataset, showcasing the effectiveness of small-scaled models when handling limited datasets.
IV-B9 Impact of Training Datasets
@@ V Conclusion
How do different YOLO versions perform on datasets with diverse object characteristics?
What are the specific strengths and limitations of each YOLO version, and how can these insights inform the selection of suitable algorithms for various applications?
In particular, the YOLO11 family emerged as the most consistent, with YOLO11m striking an optimal balance between accuracy, efficiency, and model size. While YOLOv10 delivered slightly lower accuracy than YOLO11, it excelled in speed and efficiency, making it a strong choice for applications requiring efficiency and fast processing. Additionally, YOLOv9 performed well overall and particularly stood out in smaller datasets. These findings provide valuable insights for industry and academia, guiding the selection of the most suitable YOLO algorithms and informing future developments and enhancements.
YOLOv12, despite being the newest addition to the YOLO family, delivered underwhelming performance. Its ambitious architectural changes, including the Area Attention Module and R-ELAN, introduced complexity without translating into clear advantages. The model struggled to balance accuracy and speed, highlighting the challenges of integrating advanced attention mechanisms effectively within the YOLO framework.
While the evaluated algorithms show promising performance, there is room for improvement. Future research should focus on optimizing YOLOv10’s accuracy while maintaining its speed and efficiency. Additionally, simplifying YOLOv12’s complex attention mechanisms could enhance its speed without compromising accuracy. Further architectural advancements could lead to even more efficient and accurate YOLO models. Our future work will address these gaps and explore enhancements to maximize overall efficiency and real-world applicability.